Technology stocks have an expected rate of return of 15%. MS is a computer company with a stock price of $100 per share. The company is going to pay a dividend of $5 per share at the end of the year.

a) According to the Gordon growth model, what is the market's forecast of the annual dividend growth rate for MS?

b) The market's dividend growth forecasts for MS are suddenly revised downward to 5% per year. What happens to the price of MS? What happens to the dividend-price ratio?
Answer #1

1.
Price=Expected Dividend/(expected return-growth rate)
=>100=5/(15%-g)
=>g=15%-5/100
=>g=10%

2.
Price=Expected Dividend/(expected return-growth rate)=5/(15%-5%)=50

Dividend Price ratio=5/50=10%